# Effectively Leveraging Momentum Terms in Stochastic Line Search Frameworks for Fast Optimization of Finite-Sum Problems

### v3

@@ 5 The Algorithmic Framework
  • employ the damping strategy from [fan2023];
  • compute the direction according to the strategy proposed in [lapucci2024globally], i.e.,
@@
    In practical implementations, the algorithm will often be designed to work in a random-reshuffling fashion [gurbuzbalaban2021random]. However, for the sake of simplicity and generality, we describe it in a “pure” stochastic setup.
@@
    We do not specify the particular way of choosing the tentative step size $\alpha_{0}^{k}$ at line 14. There are various options that might work both theoretically and computationally. Of course, a fixed value could be used for all iterations; however, we might prefer to define it adaptively. For this latter approach, a heuristic method, such as the one proposed by Vaswani et al. [vaswani2019], could be considered, or we could use the generalized Stochastic Polyak step size introduced by Wang et al. [wang2023].
@@ 5.1.2 Properties of the Search Direction
Together with the assumption on gradient estimate unbiasedness, largely discussed in the previous subsection, we need to state another important assumption for properly analyzing the algorithm: the search directions $d_{k}$ shall in fact be somehow tied to the true gradient $\nabla f(x^{k})$. As pointed out in well-established literature (see, e.g., [bottou2018, sec. 4.1]), the sequence of search directions $\{d_{k}\}$ shall be such that, for all $k$,
$\mathbb{E}_{k}[\|d_{k}\|]\leq C_{1}\|\nabla f(x^{k})\|,\quad\text{ and }\quad\mathbb{E}_{k}[d_{k}]^{T}\nabla f(x^{k})\leq-\|\nabla f(x^{k})\|^{2}.$
Following the discussion in [lapucci2024convergenceconditionsstochasticline], this can be ensured if, given constants $c_{1}$, $c_{2}$ and $c_{3}$, the entire sequence $\{d_{k}\}$ satisfies the following set of conditions:
@@
In order to ensure that condition (15) is also satisfied by the direction resulting from the safeguard condition, we need to keep parameter $\beta_{k}$ bounded; this is also not particularly restrictive, as it can be enforced by a clipping operation at the update time; in other words, we can simply set $\beta_{k+1}=\min\{\hat{\beta}_{k+1},\bar{\beta}\}$, where $\hat{\beta}_{k+1}$ is the value resulting from the employed CG update rule. With this precaution, we only have to rely on an assumption on the variance of $\beta_{k+1}$, which appears reasonable as argued in [lapucci2024convergenceconditionsstochasticline, Remark 2].
@@ 5.1.3 Convergence Results
We are now ready to present the asymptotic convergence analysis for Algorithm 1. In particular, we are showing how the method proposed in this paper possesses analogous properties as the other well-known stochastic line-search based methods from the literature [vaswani2019fast, vaswani2019, loizou2021, galli2023, lapucci2024convergenceconditionsstochasticline]. Specifically, it is possible to guarantee fast convergence of the algorithmic framework under interpolation and PL assumptions [ma2018, mishkin2020interpolation, polyak1987introduction].
We recall that the minimizer interpolation property [mishkin2020interpolation, Def. 3] states that
@@
We shall also recall that (see, e.g., [lapucci2024convergenceconditionsstochasticline, Sec. 3]), $L$-smoothness, minimizers interpolation and the PL condition imply the Strong Growth Condition to hold for some $\rho>0$, i.e.,
@@
The main convergence theorem is stated hereafter. The result is recovered by [lapucci2024convergenceconditionsstochasticline, Th. 5.3].
@@
The sequence of directions $\{d_{k}\}$ satisfies the conditions (13)-(14) with constants $c_{1}$ and $c_{2}$ as a result of steps 11-14. By Proposition 5.2 and the instructions of the algorithm, it also satisfies condition (15) with $c_{3}$ defined as in the statement of the theorem. All the assumptions of [lapucci2024convergenceconditionsstochasticline, Theorem 5.3] are thus satisfied by the ones here, and we consequently get the thesis.
@@ 6 Computational Experiments
In Figure LABEL:fig:CG_rule_comparison we reported the behavior of MBCG-DP when comparing the CG rules of Fletcher-Reeves (FR), Polyak-Polak-Ribiére (PPR) and Hestenes-Stiefel (HS). In this experiment the generalized SPS (7) with $c=1$ is used to select $\alpha_{0}^{k}$ and the momentum clipping approach (contraction factor 0.5) from [fan2023] is employed in case $d_{k}$ is not a descent direction. The Fletcher-Reeves rule clearly resulted the most effective one in both tasks and for both batch sizes.
@@
In Figure LABEL:fig:CG_tentative_step_comparison we show the behavior of MBCG-DP when using different approaches for selecting the tentative step $\alpha_{0}^{k}$ for the line search: a constant value, the heuristic employed in [vaswani2019, fan2023] and the generalized SPS (7) with $c=1$ were tested. The Fletcher-Reeves rule is used in this experiment; in case $d_{k}$ is not a descent direction the repeated momentum clipping approach is used, again with contraction factor of $0.5$. The generalized SPS proved to be the most effective choice.
Finally, in Figure LABEL:fig:CG_recover_direction we observe the impact of different recovery options to retrieve a descent direction, namely: switching to the simple (negative) stochastic gradient direction, recomputing the direction according to the subspace optimization strategy proposed in [lapucci2024globally], inverting the direction and taking $d_{k}$ and adopting the repeated momentum clipping approach (contraction factor of $0.5$) from [fan2023]. In this experiment, the Fletcher-Reeves rule is employed in combination with the generalized SPS initial step size. The clipping strategy resulted the preferable choice especially in nonconvex problems.
@@
The other algorithms included in the following experiments are: vanilla SGD with momentum, Adam [kingma2015], SLS [vaswani2019], PoNoS [galli2023] and MSL_SGDM [fan2023]. The benchmark of test problems include three convex instances, consisting in the training of an RBF-kernel classifier on mushrooms, ijcnn and rcv1 datasets444datasets available at csie.ntu.edu.tw/ cjlin/libsvmtools/datasets. Moreover, three nonconvex learning tasks are taken into account: once again a multi-layer perceptron with 1000 hidden units and ReLU activations to be trained on MNIST, a three-layered convolutional architecture with 32 channels each, followed by a fully-connected layer of 512 units to be trained on FashionMNIST [xiao2017fashionMNIST] dataset555available at github.com/zalandoresearch/fashion-mnist and a ResNet18 architecture to be trained on CIFAR10 [krizhevsky2009learning] dataset666available at cs.toronto.edu/ kriz/cifar.html. For the latter network, we employed the normalization-free setup proposed in [civitelli2023]. Batch sizes of 128 and 512 were again considered for all problems.

### v4

@@ 5 The Algorithmic Framework
  • employ the damping strategy from [25];
  • compute the direction according to the strategy proposed in [29], i.e.,
@@
    In practical implementations, the algorithm will often be designed to work in a random-reshuffling fashion [39]. However, for the sake of simplicity and generality, we describe it in a “pure” stochastic setup.
@@
    We do not specify the particular way of choosing the tentative step size $\alpha_{0}^{k}$ at line 14. There are various options that might work both theoretically and computationally. Of course, a fixed value could be used for all iterations; however, we might prefer to define it adaptively. For this latter approach, a heuristic method, such as the one proposed by Vaswani et al. [17], could be considered, or we could use the generalized Stochastic Polyak step size introduced by Wang et al. [24].
@@ 5.1.2 Properties of the Search Direction
Together with the assumption on gradient estimate unbiasedness, largely discussed in the previous subsection, we need to state another important assumption for properly analyzing the algorithm: the search directions $d_{k}$ shall in fact be somehow tied to the true gradient $\nabla f(x^{k})$. As pointed out in well-established literature (see, e.g., [4, sec. 4.1]), the sequence of search directions $\{d_{k}\}$ shall be such that, for all $k$,
$\mathbb{E}_{k}[\|d_{k}\|]\leq C_{1}\|\nabla f(x^{k})\|,\quad\text{ and }\quad\mathbb{E}_{k}[d_{k}]^{T}\nabla f(x^{k})\leq-C_{2}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\|\nabla f(x^{k})\|^{2}.$
Following the discussion in [30], this can be ensured if, given constants $c_{1}$, $c_{2}$ and $c_{3}$, the entire sequence $\{d_{k}\}$ satisfies the following set of conditions:
@@
In order to ensure that condition (15) is also satisfied by the direction resulting from the safeguard condition, we need to keep parameter $\beta_{k}$ bounded; this is also not particularly restrictive, as it can be enforced by a clipping operation at the update time; in other words, we can simply set $\beta_{k+1}=\min\{\hat{\beta}_{k+1},\bar{\beta}\}$, where $\hat{\beta}_{k+1}$ is the value resulting from the employed CG update rule. With this precaution, we only have to rely on an assumption on the variance of $\beta_{k+1}$, which appears reasonable as argued in [30, Remark 2].
@@ 5.1.3 Convergence Results
We are now ready to present the asymptotic convergence analysis for Algorithm 1. In particular, we are showing how the method proposed in this paper possesses analogous properties as the other well-known stochastic line-search based methods from the literature [42, 17, 16, 18, 30]. Specifically, it is possible to guarantee fast convergence of the algorithmic framework under interpolation and PL assumptions [13, 43, 34].
We recall that the minimizer interpolation property [43, Def. 3] states that
@@
We shall also recall that (see, e.g., [30, Sec. 3]), $L$-smoothness, minimizers interpolation and the PL condition imply the Strong Growth Condition to hold for some $\rho>0$, i.e.,
@@
The main convergence theorem is stated hereafter. The result is recovered by [30, Th. 5.3].
@@
The sequence of directions $\{d_{k}\}$ satisfies the conditions (13)-(14) with constants $c_{1}$ and $c_{2}$ as a result of steps 11-14. By Proposition 5.2 and the instructions of the algorithm, it also satisfies condition (15) with $c_{3}$ defined as in the statement of the theorem. All the assumptions of [30, Theorem 5.3] are thus satisfied by the ones here, and we consequently get the thesis.
@@ 6 Computational Experiments
In Figure LABEL:fig:CG_rule_comparison we reported the behavior of MBCG-DP when comparing the CG rules of Fletcher-Reeves (FR), Polyak-Polak-Ribiére (PPR) and Hestenes-Stiefel (HS). In this experiment the generalized SPS (7) with $c=1$ is used to select $\alpha_{0}^{k}$ and the momentum clipping approach (contraction factor 0.5) from [25] is employed in case $d_{k}$ is not a descent direction. The Fletcher-Reeves rule clearly resulted the most effective one in both tasks and for both batch sizes.
@@
In Figure LABEL:fig:CG_tentative_step_comparison we show the behavior of MBCG-DP when using different approaches for selecting the tentative step $\alpha_{0}^{k}$ for the line search: a constant value, the heuristic employed in [17, 25] and the generalized SPS (7) with $c=1$ were tested. The Fletcher-Reeves rule is used in this experiment; in case $d_{k}$ is not a descent direction the repeated momentum clipping approach is used, again with contraction factor of $0.5$. The generalized SPS proved to be the most effective choice.
Finally, in Figure LABEL:fig:CG_recover_direction we observe the impact of different recovery options to retrieve a descent direction, namely: switching to the simple (negative) stochastic gradient direction, recomputing the direction according to the subspace optimization strategy proposed in [29], inverting the direction and taking $d_{k}$ and adopting the repeated momentum clipping approach (contraction factor of $0.5$) from [25]. In this experiment, the Fletcher-Reeves rule is employed in combination with the generalized SPS initial step size. The clipping strategy resulted the preferable choice especially in nonconvex problems.
@@
The other algorithms included in the following experiments are: vanilla SGD with momentum, Adam [8], SLS [17], PoNoS [18] and MSL_SGDM [25]. The benchmark of test problems include three convex instances, consisting in the training of an RBF-kernel classifier on mushrooms, ijcnn and rcv1 datasets444datasets available at csie.ntu.edu.tw/ cjlin/libsvmtools/datasets. Moreover, three nonconvex learning tasks are taken into account: once again a multi-layer perceptron with 1000 hidden units and ReLU activations to be trained on MNIST, a three-layered convolutional architecture with 32 channels each, followed by a fully-connected layer of 512 units to be trained on FashionMNIST [44] dataset555available at github.com/zalandoresearch/fashion-mnist and a ResNet18 architecture to be trained on CIFAR10 [45] dataset666available at cs.toronto.edu/ kriz/cifar.html. For the latter network, we employed the normalization-free setup proposed in [46]. Batch sizes of 128 and 512 were again considered for all problems.